OpenStudy (olivya999):

Confirm that f and g are inverses by showing that f(g(x)) = x and g(f(x)) = x.

OpenStudy (olivya999):

\[f(x) = \frac{ 4}{ x } \] and \[g(x)= \frac{ 4 }{ x }\]

zepdrix (zepdrix):

\[\large\rm f(\color{orangered}{x})=\frac{4}{\color{orangered}{x}}\]Ok let's try one of the directions, let's plug g into our f,\[\large\rm f(\color{orangered}{g(x)})=\frac{4}{\color{orangered}{g(x)}}\]Replacing g(x) with 4/x in our function gives us,\[\large\rm f(\color{orangered}{g(x)})=\frac{4}{\color{orangered}{\frac4x}}\]

zepdrix (zepdrix):

Remember what to do when dividing by a fraction? Have you learned `Keep Change Flip` or something similar perhaps?

OpenStudy (olivya999):

yeah

OpenStudy (olivya999):

@zepdrix

zepdrix (zepdrix):

\(\large\rm \dfrac{a}{\frac{b}{c}}=a\cdot\frac{c}{b}\) `Keep` the numerator the same, `Change` the operation from division to multiplication, `Flip` the bottom fraction.

OpenStudy (olivya999):

so it would be \[4 \times \frac{ x }{ 4 }\]

zepdrix (zepdrix):

Yes, and then deal with the 4's :) Cancellation, yes?

OpenStudy (olivya999):

yes so \[1\times x\]

OpenStudy (olivya999):

@zepdrix

zepdrix (zepdrix):

Ya so it looks like f(g(x)) = x Good.

OpenStudy (olivya999):

thank you @zepdrix

zepdrix (zepdrix):

Can you try the other direction on your own? :) It should be very similar.

zepdrix (zepdrix):

Oh, actually, it's exactly the same lol